OpenStudy (anonymous):

MEDAL AND FAN! A researcher in North America discovered a fossil that contains 78% of its original amount of c-14. What's the age of the fossil rounded to the nearest year?

OpenStudy (anonymous):

google it that is out of my league

OpenStudy (anonymous):

@adamaero help?

OpenStudy (adamaero):

Do you think you need to know the half-life of c-14?

OpenStudy (anonymous):

probably because that's what this unit is over, but it doesn't look like it'd be done the same way as the other..

OpenStudy (adamaero):

I'm not sure, so try to look at the answers in the back of the book of the similar problems around that one or at the examples in the beginning of the chapter. If you still don't get it, go to the math homepage and ask ppl that are near the top w/ @ like you did with me.

OpenStudy (adamaero):

Find N not first

OpenStudy (anonymous):

Alright, thanks I'll go back through the tutorial and see if I can figure it out.

OpenStudy (adamaero):

you want to find all the terms before you find t (the time in years--or is it some other form?)

OpenStudy (anonymous):

pretty sure time in years

OpenStudy (wolf1728):

Carbon 14 half life = 5,730 years Rather than rely on that nasty formula you posted here's one that's easier to understand: elapsed time = half life * log (beginning amount/ending amount) / log 2 elapsed time = 5,730 * log (100 / 78) / log 2 elapsed time = 5,730 * log (1.2820512821) / 0.30102999566 elapsed time = 5,730 * (0.10790539733 / 0.30102999566) elapsed time = 2,054 years. Half life calculator and formulas http://www.1728.org/halflife.htm

OpenStudy (adamaero):

ya, but that is the format of the question on a test, the same equation will be giving and the half life of carbon-bla will not be given

OpenStudy (wolf1728):

adamaero You cannot do this problem without the half life of carbon 14 Also the formula given will have to be reworked mathematically because as it is given, it is solved for present amount and NOT solved for time.